OpenStudy (anonymous):

PLEASE HELP! Give examples of 3 different polynomial identities

OpenStudy (ajanijones):

(a+b) 2 = a 2 + 2ab + b 2 (a+b)(c+d) = ac + ad + bc + bd a 2 - b 2 = (a+b)(a-b) (Difference of squares) a 3 (+-) b 3 = (a (+-) b)(a 2(-+) ab + b 2) (Sum and Difference of Cubes) x 2 + (a+b)x + AB = (x + a)(x + b) if ax 2 + bx + c = 0 then x = ( -b (+-)sqrt(b 2 - 4ac) ) / 2a (Quadratic Formula)

OpenStudy (anonymous):

I can help you

OpenStudy (ajanijones):

Hope I helped out :)

OpenStudy (ajanijones):

@cobra-strikes id you get it?

OpenStudy (anonymous):

i looked at those problems and i think my brain just died xD @ajanijones

OpenStudy (ajanijones):

xd its pretty easy but it takes time, once you get the hang of it it will start to flow like water in a stream.

OpenStudy (anonymous):

sfgkjrihbgdgfehgi ghfh gejghkql qtthreghdlkjs ehshwh xhk,sxkzsnma kd

OpenStudy (anonymous):

#gogirl